# TEOChat: A Large Vision-Language Assistant for Temporal Earth Observation Data

### v1

@@ 4.2 TEOChat Training Details
Following prior work (Kuckreja et al., 2023), to retain the strong capabilities of the pre-trained image encoder and LLM, and to minimize memory usage during training, we freeze the vision encoder and projector weights then fine-tune the LLM using Low-Rank Adaptation (LoRA) (Hu et al., 2021). To further reduce memory footprint, we use 8-bit quantization of the base LLM weights. These strategies allow us to train the large multimodal model on temporal sequences of up to 8 images using an NVIDIA A4000 GPU (16GB VRAM). We provide additional training details in the Appendix.
@@ 5 Experimental Results
Table 1: Temporal task performance of TEOChat compared to other VLMs and specialist approaches. We bold the best generalist result on each task. TEOChat outperforms a VLM instruction-tuned on natural images and videos (Video-LLaVA (Lin et al., 2023)) as well as a VLM instruction-tuned on single EO images (GeoChat (Kuckreja et al., 2023)). TEOChat also performs comparably to or outperforms a specialist model trained to perform specific tasks (fMoW (Cong et al., 2022), xBD (Gupta et al., 2019), S2Looking (Shen et al., 2021), QFabric (Verma et al., 2021)). TEOChat-T was only trained on temporal tasks.
We evaluate TEOChat in a variety of settings. First, we find that TEOChat demonstrates impressive temporal reasoning capabilities, outperforming two strong VLMs, one for videos (Video-LLaVA) and one for single EO images (GeoChat), and even rivals or outperforms specialist approaches (Section 5.1). Second, we test model design decisions in experimental ablations, and find our design of TEOChat is superior to many others (Section 5.2). Third, we demonstrate that TEOChat achieves impressive zero-shot generalization to new temporal datasets not in TEOChatlas  (Section 5.3). Fourth, we find that TEOChat outperforms two strong proprietary foundation models trained to model sequences of images (Section 5.4). Fifth, we observe that TEOChat possesses strong single image capabilities, outperforming GeoChat on average across zero-shot scene classification and visual question answering (Section 5.5).
@@ 5.1 Comparison to other VLMs and specialist approaches
Baselines and Evaluation Metrics  We compare TEOChat to Video-LLaVA (Lin et al., 2023), GeoChat (Kuckreja et al., 2023), and specialist models (Gupta et al., 2019; Shen et al., 2021; Verma et al., 2021). To obtain GeoChat’s predictions on the temporal tasks, we perform temporal post-processing of its single image predictions. For example, for building change detection, we obtain building location predictions in each image separately, then compute a per-pixel difference to identify the changed buildings over time. We describe all temporal post-processing strategies in the Appendix. For the specialist models, we use a strong self-supervised EO model adapted to fMoW (SatMAE (Cong et al., 2022)) and strong models trained on each individual dataset for the change detection tasks (a modified UNet (Ronneberger et al., 2015) from Gupta et al. (2019), FC-Siam-Diff (Daudt et al., 2018) from Shen et al. (2021), and a modified UNet from Verma et al. (2021)).
We use task-specific evaluation metrics. For previously existing tasks, we follow the metrics reported in the original works. For new tasks, we use accuracy. See Appendix for a full description of metrics, and see Figure 4 for qualitative examples of TEOChat responses on the temporal tasks.
Temporal Scene Classification  TEOChat achieves impressive performance on TSC with the fMoW RGB (75.1%) and Sentinel (45.5%) validation sets, outperforming both Video-LLaVA (16.6%, 4.9%) and GeoChat (59.2%, 26.3%) (Table 1). We also compare TEOChat to linear probing a specialist encoder that was pre-trained with self-supervision on fMoW images (SatMAE). TEOChat obtains considerably higher accuracy than SatMAE on fMoW RGB (+9.2%), and emphasize the TEOChat is a generalist which is not pre-trained on EO imagery. We also note that, impressively, TEOChat performs comparably to SatMAE+Stack fully fine-tuned on fMoW RGB (-0.8%).
Change Detection and Spatial Referring Expression  TEOChat outperforms Video-LLaVA and GeoChat by a substantial margin on all CD tasks (Table 1). However, on xBD building localization, the model underperforms the specialist approach by a large margin. Qualitatively, the model does moderately well at identifying the building locations but struggles to precisely localize their boundaries. This may be partially due to representing object locations using natural language tokens, and partially due to using axis-aligned boxes rather than general polygons. TEOChat’s change detection performance on S2Looking is higher than the specialist model’s performance (+7.1 F1). On building damage classification, TEOChat outperforms the specialist substantially (+23.5 F1). On SRE, TEOChat achieves much higher performance than both Video-LLaVA (+19.6 F1 on xBD, +30.6 F1 on S2Looking) and GeoChat (+14.5 F1 on xBD, +27.8 F1 on S2Looking).
@@ 5.2 Model ablations
Temporal task vs. joint single-temporal training. We investigate the effect of joint single image and temporal task training by training a model on only the temporal tasks which we call TEOChat-T. TEOChat-T underperforms TEOChat on all but two tasks, matching TEOChat on fMoW Sentinel and slightly outperforming it on S2Looking QA (Table 1). This suggests that joint training on single image tasks also largely benefits temporal reasoning capabilities.
Model design We test various design decisions of TEOChat by comparing their performance on the canonical temporal tasks in TEOChatlas, namely fMoW RGB, xBD damage classification, S2Looking change detection, and QFabric RQA with 2 and 5 images. Specifically, we test the impact of using a LLaVA initialization instead of Video-LLaVA, using an image encoder initialization from strong remote sensing pre-trained weights (SkyScript (Wang et al., 2024b)) instead of CLIP, fine-tuning versus freezing the projector layer, and including image references in the prompt. We note that we couple the LLM and projector initialization as they were pretrained together. Furthermore, to fairly compare to a new image encoder initialization, we also test the impact of using a fine-tuned projector with the alternate initialization to allow it to adapt to the new image encoder outputs. Finally, we measure the performance improvement from training on TEOChatlas (without using image references) and the effect of training for longer (7k iters [1 epoch] vs 14k [2 epochs]).
Our design of TEOChat outperforms all tested designs on the five canonical temporal tasks (Table 2). Fine-tuning on TEOChatlas leads to large performance improvements across the tasks and including image references leads to further improvements. We further measure the impact of including image identifiers in the test-time prompt input to TEOChat for temporal referring tasks (Table 9), and find that the identifiers lead to substantial performance improvements on both tasks. Using LLaVA weights for the LLM and projector initialization underperforms VideoLLaVA across all tasks which we attribute to the fact that Video-LLaVA is a stronger model and was trained with natural video data. The use of remote sensing pre-trained image encoder also underperforms a CLIP-pretrained image encoder even when allowing projector fine-tuning, which we believe could be partially because the LLM weights were pre-trained with a frozen CLIP-initialized image encoder during multi-modal instruction-tuning on natural video data. We also find that fine-tuning outperforms freezing the projector, and training the model for longer leads to further performance improvements on all tasks, with especially large improvements on fMoW RGB and xBD damage classification.
@@ 5.3 Zero-shot performance on change detection tasks
We assess TEOChat’s performance on two new temporal remote sensing datasets, namely change detection using ABCD (Fujita et al., 2017) and change visual question answering using CDVQA (Yuan et al., 2022). TEOChat demonstrates impressive generalization to thse new datasets, outperforming Video-LLaVA (+35.6 on ABCD, +17.4 on CDVQA) and approaching specialist performance.
5.4 Comparison to proprietary foundation models
We compare TEOChat to GPT-4o and Gemini 1.5 Pro with three demonstrating examples on xBD damage classification and S2Looking change detection (Table 4). TEOChat performs better than both proprietary foundation models with in-context learning on xBD damage classification (+11.7 F1 compared to GPT-4o, +12.8 F1 compared to Gemini 1.5 Pro) and S2Looking change detection (+14.2 F1 compared to GPT-4o, +17.1 F1 compared to Gemini 1.5 Pro). We could not evaluate the models on fMoW or QFabric (which are larger both in size and sequence length) due to cost.
@@ 5.4 Comparison to proprietary foundation models
Table 3: Zero-shot results on ABCD (building CD) and CDVQA (change QA).
@@
Table 4: Comparison to proprietary foundation models. GPT-4o and Gemini 1.5 Pro use in-context learning with 3 examples.
@@
Table 5: Single image accuracies for scene classification (AID, UCMerced) and visual question answering (LRBEN, HRBEN). TEOChat has strong single image capabilities, outperforming both VLMs on average.
@@ 5.5 Single image performance
TEOChat achieves better performance than both models on all but one single image task (Table 5). It outperforms GeoChat on both LULC classification tasks, notably by a large margin on AID (+8.9), a much larger-scale dataset than UCMerced. TEOChat also outperforms GeoChat on the low resolution VQA tasks and the high resolution presence task, while slightly underperforming on the high resolution comparison task. Notably, GeoChat uses higher resolution images than TEOChat (504 vs. 224). TEOChat attains the highest average performance by a considerable margin (+3.5), suggesting it has stronger single image capabilities on top of the new temporal reasoning capabilities. We provide examples of TEOChat single image responses in Figure 10.
@@ 6 Conclusion
In this work, we propose TEOChat, the first vision and language assistant that can engage in conversation about temporal EO imagery. We introduce a novel dataset called TEOChatlas to train TEOChat on a diverse set of spatial and temporal instruction-following tasks. Our experiments suggest that TEOChat exhibits strong performance, outperforming previous VLMs substantially, and even competing with or outperforming specialist models. TEOChat also demonstrates impressive zero-shot generalization to two change detection datasets and outperforms strong proprietary models which can handle sequences of images.
@@
Dataset Size Temporal Detection

### v2

@@ 4.2 TEOChat Training Details
Following prior work (Kuckreja et al., 2023), to retain the strong capabilities of the pre-trained image encoder and LLM, and to minimize memory usage during training, we freeze the vision encoder then fine-tune the LLM using Low-Rank Adaptation (LoRA) (Hu et al., 2021). To further reduce memory footprint, we use 8-bit quantization of the base LLM weights. These strategies allow us to train the large multimodal model on temporal sequences of up to 8 images using an NVIDIA A4000 GPU (16GB VRAM). We provide additional training details in Appendix Section F.
@@ 5 Experimental Results
Table 1: Temporal task performance of TEOChat compared to other VLMs and specialists. We bold the best generalist result on each task. TEOChat outperforms a VLM instruction-tuned on natural images and videos (Video-LLaVA) as well as a VLM instruction-tuned on single EO images (GeoChat). We note Video-LLaVA and GeoChat are not trained on these tasks. TEOChat also performs comparably to or outperforms several specialist models trained to perform specific tasks. TEOChat-T was only trained on temporal tasks.
We evaluate TEOChat in a variety of settings. First, we find TEOChat demonstrates impressive temporal reasoning capabilities, outperforming two strong VLMs, one for videos (Video-LLaVA) and one for single EO images (GeoChat), and even rivals or outperforms several specialist approaches (Section 5.1). Second, through experimental ablations we show TEOChat’s design is superior to many others (Section 5.2). Third, TEOChat achieves impressive zero-shot generalization to new temporal datasets not in TEOChatlas  (Section 5.3). Fourth, we find that TEOChat outperforms two strong proprietary VLMs trained on sequences of images (Section 5.4). Fifth, we show TEOChat possesses strong single image capabilities, outperforming GeoChat on average across zero-shot scene classification, VQA, and captioning (Section 5.5). Sixth, we find TEOChat exhibits interesting instances of generalization to tasks not explicitly represented in the training set (Section 5.6).
@@ 5.1 Comparison to other VLMs and specialist approaches
Baselines and Evaluation Metrics  We compare TEOChat to Video-LLaVA (Lin et al., 2023), GeoChat (Kuckreja et al., 2023), and specialist models. To obtain GeoChat’s predictions on the temporal tasks, we perform temporal post-processing of its single image outputs (Appendix Section G). For the specialists, we use a strong self-supervised EO model adapted to fMoW (SatMAE (Cong et al., 2022)) and strong models trained on each individual dataset for the change detection tasks (a modified UNet (Ronneberger et al., 2015) from Gupta et al. (2019), FC-Siam-Diff (Daudt et al., 2018) from Shen et al. (2021), and a modified UNet from Verma et al. (2021)). We evaluate using task-specific metrics described in detail in Appendix Section H.
Temporal Scene Classification  TEOChat achieves impressive performance on TSC with the fMoW RGB (75.1%) and Sentinel (45.5%) validation sets, outperforming both Video-LLaVA (16.6%, 4.9%) and GeoChat (59.2%, 26.3%) (Table 1). We also compare TEOChat to linear probing a specialist encoder that was pre-trained with self-supervision on fMoW images (SatMAE). TEOChat obtains considerably higher accuracy than SatMAE on fMoW RGB (+9.2%), and emphasize the TEOChat is a generalist which is not pre-trained on EO imagery. We also note that, impressively, TEOChat performs comparably to SatMAE+Stack (-0.8%) and is approaching state-of-the-art (SOTA) specialist performance on fMoW RGB (-4.1%) (Khanna et al., 2024a).
Change Detection and Spatial Referring Expression  TEOChat outperforms Video-LLaVA and GeoChat by a substantial margin on all CD tasks (Table 1). However, on xBD localization, it underperforms the specialist by a large margin. TEOChat’s CD performance on S2Looking is higher than the specialist model’s performance (+8.0 F1). However, we note that SOTA performance on S2Looking CD (Li et al., 2024b) and xBD localization Chen et al. (2024a) is much higher than TEOChat (+31.3 F1, +48.5 F1), which we believe is largely because these bitemporal change detection models output segmentation masks, whereas TEOChat is constrained to output boxes. Qualitatively, TEOChat does moderately well at identifying building locations but struggles to precisely localize boundaries and localize large numbers of buildings. This may be partially due to representing object locations using natural language tokens, and partially due to using axis-aligned boxes rather than general polygons. On damage classification, TEOChat outperforms the specialist substantially (+23.5 F1). On SRE, TEOChat achieves much higher performance than Video-LLaVA (+21.4 F1 on xBD, +31.2 F1 on S2Looking) and GeoChat (+16.4 F1 on xBD, +28.4 F1 on S2Looking).
@@ 5.2 Model ablations
Temporal vs. joint single-temporal training. We investigate the effect of joint single image and temporal training by training a model on only the temporal tasks (TEOChat-T). TEOChat-T underperforms TEOChat on all but two tasks, matching it on fMoW Sentinel and slightly outperforming it on S2Looking QA (Table 1). This suggests joint training also largely benefits temporal capabilities.
Model design We test various design decisions of TEOChat. Specifically, we measure the impact of using a LLaVA initialization instead of Video-LLaVA, using an image encoder initialization from strong remote sensing pre-trained weights (SkyScript (Wang et al., 2024d)) instead of CLIP, fine-tuning versus freezing the projector layer, and including image references in the prompt. We note that we couple the LLM and projector initialization as they were pretrained together. Furthermore, to fairly compare to a new image encoder initialization, we also test the impact of using a fine-tuned projector with the alternate initialization to allow it to adapt to the new image encoder outputs. Finally, we measure the performance improvement from training on TEOChatlas (without using image references) and the effect of training for longer (7k iters [1 epoch] vs 14k [2 epochs]).
Our design of TEOChat outperforms all tested designs on the five canonical temporal tasks (Table 2). Fine-tuning on TEOChatlas leads to large performance improvements across the tasks and including image references leads to further improvements. Interestingly, image references help even in the bitemporal case, which is likely because the ‘Image 2‘ reference helps delineate the visual tokens between the two images, as it explains most of the improvement from including references on the canonical bitemporal tasks (Table 9). The benefits of image references hold at test-time as well, as their exclusion from the test-time prompt leads to substantial performance drops on both tasks (Table 10). Using LLaVA weights for the LLM and projector initialization underperforms VideoLLaVA across all tasks which we attribute to the fact that Video-LLaVA is a stronger model and was trained with natural video data. The use of remote sensing pre-trained image encoder also underperforms a CLIP-pretrained image encoder even with projector fine-tuning, which we believe is partially because the LLM weights were pre-trained with a frozen CLIP-initialized image encoder during multi-modal instruction-tuning on natural video. We also find that fine-tuning outperforms freezing the projector, and training the model for longer leads to further performance improvements on all tasks, with especially large improvements on fMoW RGB and xBD damage classification.
@@ 5.3 Zero-shot performance on change detection tasks
We assess TEOChat on two new temporal remote sensing datasets, namely change detection using ABCD (Fujita et al., 2017) and change visual question answering using CDVQA (Yuan et al., 2022) (Table 4). TEOChat demonstrates impressive generalization to these datasets, outperforming Video-LLaVA (+35.6 on ABCD, +17.4 on CDVQA) and approaching specialist performance.
5.4 Comparison to other multi-image VLMs
We compare TEOChat to open-source multi-image VLMs InternVL2-4B (Chen et al., 2024b) and Qwen2VL-7B-Instruct (Wang et al., 2024b) as well as GPT-4o and Gemini 1.5 Pro with three demonstrating examples on xBD damage classification and S2Looking change detection (Table 4). We focus this analysis on these datasets as we could not evaluate the proprietary models on fMoW or QFabric (larger both in size and sequence length) due to cost. TEOChat outperforms both Qwen2-VL and InternVL2 by a large margin on xBD damage classification (+38.2 F1, +32.7 F1 respectively) and S2Looking change detection (+25.8 F1, +23.7 F1 respectively). TEOChat also performs better than both proprietary models on xBD damage classification (+11.7 F1 compared to GPT-4o, +12.8 F1 compared to Gemini 1.5 Pro) and S2Looking change detection (+14.2 F1 compared to GPT-4o, +17.1 F1 compared to Gemini 1.5 Pro). Qualitatively, the proprietary models tend to underpredict change in S2Looking and underestimate building damage in xBD, whereas the open-source models produce highly inaccurate boxes and classify almost all examples as ‘No Damage’.
@@ 5.4 Comparison to other multi-image VLMs
Table 3: Zero-shot ABCD (building CD) and CDVQA (change QA) results.
@@
Table 4: Comparison to other multi-image VLMs. GPT-4o and Gemini 1.5 Pro use three in-context examples.
@@
RSVQA (Lobry et al., 2020) - - 87.5 81.5 90.0 90.4 88.2 -
RSGPT (Hu et al., 2023) - - 91.2 91.7 94.0 90.9 90.0 -
LHRS-Bot (Muhtar et al., 2024) 91.3 - 88.5 90.0 89.1 92.6 92.5 -
VHM (Pang et al., 2024) 91.7 - 91.1 92.0 95.0 64.0 83.5 -
@@
Table 5: Single image accuracies for scene classification (AID, UCMerced) and VQA (LRBEN, HRBEN). We italicize results from fine-tuning on the corresponding training set. All other results are zero-shot. We bold the highest zero-shot score, except LRBEN which all models but Video-LLaVA are fine-tuned on.
@@ 5.5 Single image performance
TEOChat matches or exceeds the performance of all single image models on six of the seven tasks (Table 5). It outperforms GeoChat on both LULC classification tasks, notably by a large margin on AID (+8.9), a much larger-scale dataset than UCMerced. It underperforms LHRS-Bot and VHM on AID, likely due to the inclusion of several more LULC classification datasets in their training data. TEOChat also outperforms all models on low resolution VQA presence and comparison tasks, and attains close to the best performance on rural/urban classification. It achieves the highest zero-shot performance on high resolution (HR) presence, while slightly underperforming GeoChat and VHM on HR comparison. Notably, GeoChat and VHM use higher resolution images than TEOChat (504 and 336 vs. 224). TEOChat attains higher average performance than GeoChat by a considerable margin (+3.5), suggesting it has stronger single image capabilities on top of the new temporal ones. Finally, we measure TEOChat’s performance on three EO caption datasets. We find it outperforms GeoChat on all datasets and matches other single image EO generalists on the larger two datasets in the fine-tuning setting (Tables 11, 12, 13 and 14), and matches or outperforms GeoChat on average in the zero-shot setting (Tables 15, 16, 17 and 18). Figures 7 and 8 show example TEOChat single image responses.
5.6 TEOChat Generalization
TEOChat also exhibits interesting cases of generalization to tasks not explicitly represented in the training data (Figure 4). For example, TEOChat can compose its spatial and temporal reasoning abilities, such as identifying objects that are only regionally annotated in the single image training tasks (like airplanes and cars) within different images in the temporal sequence. It can also recognize differences between images (like seasonality) which is not explicitly reflected in the training data.
Figure 4: Qualitative examples of task generalization exhibited by TEOChat on temporal EO images.
@@ 6 Conclusion
In this work, we propose TEOChat, the first vision and language assistant that can engage in conversation about temporal EO imagery. We introduce a novel dataset called TEOChatlas to train TEOChat on a diverse set of spatial and temporal instruction-following tasks. TEOChat exhibits strong performance, outperforming previous VLMs substantially, and even competing with or outperforming several specialist models. TEOChat also demonstrates impressive zero-shot generalization to two change detection datasets and outperforms strong multi-image proprietary models .
@@
7 Reproducibility Statement
Training data, model weights, training code, and evaluation code are hosted at https://github.com/ermongroup/TEOChat. All major experimental results can be reproduced by following the steps described there.
@@ 7 Reproducibility Statement
A Related Work
ML for Temporal Earth Observation  ML methods have been commonly used to model temporal sequences of earth observation data (Moskolaï et al., 2021; Miller et al., 2024; Khanna et al., 2024b). Change detection is a widely studied earth observation task where the goal is to identify change between pairs (bitemporal) or sequences (multitemporal) of images capturing the same location over time (Chughtai et al., 2021; Bai et al., 2023; Cheng et al., 2024). Automated methods for change detection can aid important humanitarian and sustainability efforts, including planning disaster relief (Rahnemoonfar et al., 2021; Gupta et al., 2019; Ban et al., 2020), tracking urban changes (Tamilenthi & Baskaran, 2013; Fyleris et al., 2022; Verma et al., 2021), and monitoring ecological conditions (Willis, 2015; Xu et al., 2019), deforestation (Shermeyer & Haack, 2015; De Bem et al., 2020; Decuyper et al., 2022), and crops (Gim et al., 2020; Kaur et al., 2023; Wang et al., 2024a).
Commonly studied change detection ML subtasks include building change detection (Shen et al., 2021) and damage detection (Gupta et al., 2019), semantic change detection including land cover changes (Daudt et al., 2019; Yang et al., 2021; Toker et al., 2022) and land use changes (Verma et al., 2021). Many ML-based approaches have been developed to accomplish these tasks, and almost all use a Siamese architecture where the vision encoder is shared across time (Zheng et al., 2021d; b; 2022b; Liu et al., 2024c; Tian et al., 2023), which we adopt as well.
Beyond change detection, multitemporal scene classification using the Functional Map of the World (fMoW) is a common benchmark (Christie et al., 2018). Finally, while several works have designed self-supervised approaches to leverage temporal sequences of earth observation data (Ayush et al., 2021; Manas et al., 2021; Cong et al., 2022), none have developed vision-language models that can perform a wide variety of instruction-following tasks on earth observation imagery over time.
VLMs for Natural Images  There have been substantial recent advancements in the development of vision-language models (VLMs) for natural images, notably including LLaVA (Liu et al., 2024a) and its successor LLaVA-1.5 (Liu et al., 2023a). These seminal works have demonstrated the effectiveness of curating large multimodal instruction-following datasets to train high performing vision-language models that can respond to user instructions about image data using a simple model design consisting of a vision encoder to get visual representations, a projector layer to convert visual tokens into language space, and an LLM to decode the instructions and converted visual tokens into a response. Subsequent works have proposed approaches to build upon these instruction-tuned VLMs, from improving performance with mixture of experts (Lin et al., 2024), achieving similar performance with fewer parameters (Zhou et al., 2024; Zhu et al., 2024), and improve task transfer (Li et al., 2024a) to adding new capabilities like object detection (Wang et al., 2024c), semantic segmentation (Rasheed et al., 2023; Zhang et al., 2023a), image generation (Sun et al., 2023), and tool use (Liu et al., 2023b).
Temporal VLMs  Several recent works have developed VLMs that can engage in conversation about videos. VideoChatGPT curates a video instruction set to train a LLaVA-style architecture but combining both spatially pooled temporal features and temporally pooled spatial features on CLIP-encoded frames (Maaz et al., 2023). VideoLLaVA then demonstrates improved performance by training a LLaVA-style architecture with separate, pre-aligned image and video encoders on a joint instruction-following dataset of both images and videos (Lin et al., 2023). Subsequent works have primarily adopted LLaVA-style architectures as well, but include additional layers aimed to process visual tokens across time before inputting them into the LLM (Jin et al., 2023; Li et al., 2023; Zhang et al., 2023b; Liu et al., 2024b).
VLMs for EO  Efforts to develop instruction-following VLMs for EO data have been rapidly increasing (Hu et al., 2023; Kuckreja et al., 2023; Muhtar et al., 2024; Pang et al., 2024). These methods commonly curate many different single image EO datasets and convert them to instruction-following tasks, then instruction-tune a LLaVA-like model on the curated dataset. Importantly, no prior work has developed VLMs for temporal sequences for EO except for CDChat (Noman et al., 2024), a concurrent work that fine-tunes LLaVA-1.5 on two bitemporal change detection datasets. In contrast, our work builds upon the more recent VideoLLaVA architecture, enabling the handling of sequences with up to eight images. Additionally, we develop a public instruction-following dataset including many different single image tasks as well as temporal tasks from four EO datasets, broadening the scope and capabilities of the model.
Dataset Examples (Images) Temporal Detection